Quantcast

What works and what is paused · Email is limited to account activation and password resets.

Jump to content


Photo

I NEED MATH WIZES


  • Please log in to reply
10 replies to this topic

#1 Nearx

Nearx
  • 506 posts

Posted 07 March 2011 - 10:30 AM

I'm taking a quiz but its open book. I never got my book so that doesn't help much.

I came across one problem

If P(A) = .2, find P(A) (There's a line over the "A" in the last P(A) thing)

I dont want the answer to that problem, I just need to know how the heck to do it x.x
Can anyone help? Maybe give me an example problem to go off of?

Advertisement

    #2 rizler

    rizler
    • 637 posts

    Posted 07 March 2011 - 10:53 AM

    this is probability? and the second A is like this?
    _
    A?

    i think the forumula you are looking for if its probability is


    P(A) = The Number Of Ways Event A Can Occur
    The total number Of Possible Outcomes

    am studying Maths in college 1st year :)

    weve finished with probability but i think i can still remmeber some

    Edited by rizler, 07 March 2011 - 10:56 AM.


    #3 Nearx

    Nearx
    • 506 posts

    Posted 07 March 2011 - 10:58 AM

    Yup but only on the second A. Then the second question is backwards. With the lined A first.

    If P(-A)=.02 find P(A)

    - is the "line"

    This is my first college year in math too. We just finished probability and had to take a quiz on it xD
    But yes that's the formula. I just don't understand completely how to execute it

    I had to go turn in the quiz but I'd still like to know. I'm pretty sure I butchered those 3 P(A) questions x.x
    I just hope I at least did ok on the rest of it all. UGH - math T___T

    #4 rizler

    rizler
    • 637 posts

    Posted 07 March 2011 - 11:00 AM

    well to be honest to solve it your probably gonna have to right some stuff down on paper seeing as to do it on here would be hard :p

    either use a TREE Diagram
    OR
    VENN diagram (i recommend this if decimals)

    then label everything and plot in your numbers

    #5 ToxicS

    ToxicS
    • 2580 posts

    Posted 07 March 2011 - 11:04 AM

    Agh I learned this two semesters ago. Don't remember it now :(

    #6 Nearx

    Nearx
    • 506 posts

    Posted 07 March 2011 - 11:04 AM

    o.o

    <.<
    this is why I hate math xD

    There's a similar problem like this on the board:

    P(E) = 1/5 (fraction)

    Step 1: P(-E) = 1-P(E) = 1-1/5=5/5-1/5=4/5
    Step 2: P(E)/P(-E) = 1/5 / 4/5 = 1/4

    Then the answer is 1:4 or 1 to 4

    I keep thinking the P(A) one is similar to P(E).... I didn't have to use Tree or Venn diagrams for the E one so I'm a little confused =/
    The teacher actually never went over the diagrams for this section...

    #7 Boggart

    Boggart
    • Professional Napper

    • 7981 posts


    Users Awards

    Posted 07 March 2011 - 11:22 AM

    I think what it's asking is the same thing, the notation for P(E) is just different (and not usually used).

    If the probability of A occuring is noted as P (A) then propability of A NOT occuring is P(Ᾱ).

    0.2 can be represented as 2/10. So the chance of it NOT occuring is the remaining 8/10. Full work is:

    1 - 2/10 =10/10 - 2/10 = 8/10.

    No? yes? no?


    #8 Nearx

    Nearx
    • 506 posts

    Posted 07 March 2011 - 11:27 AM

    Sounds about right o-o
    xD
    I'm not sure what the answer was to that quiz but when I find out I'll look back here and compare =3
    Thanks!!!

    #9 Boggart

    Boggart
    • Professional Napper

    • 7981 posts


    Users Awards

    Posted 07 March 2011 - 11:31 AM

    I took this last year so if you need more help, feel free to ask more :)

    #10 Ladida

    Ladida
    • Night Owl 🌛

    • 2152 posts


    Users Awards

    Posted 07 March 2011 - 01:03 PM

    Stats memories :( And then you'll do the probability of something occurring/not occurring after something already occurred/did not occur. That's Baye's Theorem I think. Good luck!

    #11 Nearx

    Nearx
    • 506 posts

    Posted 07 March 2011 - 01:10 PM

    Ugh math is so frustrating x.x
    Thanks for the help guys! I can use this for my unit test =3

    Thanks Boggart! Will do :D


    1 user(s) are reading this topic

    0 members, 1 guests, 0 anonymous users